# Fast offline decoding with local message-passing automata

### v1

@@ 1.2 Comparison with prior work
The decoders most similar to ours—which served as direct inspiration for the present work—are the “field-based” CA decoders of Refs. [13, 14]. These decoders transmit information about syndrome locations in a physically natural fashion, using a scalar field $\phi$ updated according to a lattice discretization of the Poisson equation
@@
For the purposes of this work, the most significant drawback of these decoders is that they do not meet our definition of locality, with the $D_{\phi}=3$ model requiring a $(\log L)^{2}$-divergent classical communication speed (in App. D, we argue that all PDE-based models suffer from this problem). Our decoder can thus be viewed as an improvement of this construction which brings down the communication speed to an $O(1)$ constant and reduces the number of control bits used per site from $\omega((\log L)^{3})$ [13] to $\Theta(\log\log L)$.
@@
Aside from field-based decoders, the two existing types of cellular automaton decoders for the 2D toric code [^8] In this work we focus only on correcting errors that produce point-like syndromes; correcting errors with higher-dimensional syndromes (in e.g. one sector of the 3D toric code) is an easier task, and can be done using an extension of Toom’s rule, see e.g.  [16]. are Harrington’s decoder [12] and its extensions [39, 45] (see also [17, 20]), and the decoder of Ref. [15], which can be regarded as a quantum extension of Tsirelson’s automaton [46] (these decoders are designed for online decoding, but they can of course be used for offline decoding as well). Both of these decoders can loosely be thought of as eliminating anyons according to a locally-implemented RG scheme, with the different length scales involved in the RG process hard-coded into the structure of the dynamics. For the purposes of this work, drawbacks of these approaches are that they are only naturally defined on systems of size $n^{l}$ for integers $n,l$, and are numerically observed to have thresholds at rather low error rates, with the simplest formulation of Ref. [15] achieving $p_{c}\lesssim 2\%$ for offline decoding. These features make them—at least in their current form—less practical for experimental implementation and numerical study. [^9] In 2D, these decoders are also not naturally translation invariant in space or time, although this problem can be fixed formally by adding additional clock registers at each site.
@@ 1.3 Outlook
While our explicit construction is specialized to the case of $\mathbb{Z}_{2}$ topological codes on $\mathbb{Z}^{D}$, it should be straightforward to generalize the message passing architecture to arbitrary Abelian orders on general lattices (and perhaps to non-Abelian ones as well, at least phenomenologically). From an implementation perspective, it would also be valuable to quantitatively investigate different types of message passing schemes, and to systematically study how the numerical value of the threshold depends on the choice of architecture and parameters involved (such as the message transmission speed). Relatedly, it could be fruitful to study how decoding can be optimized to take advantage of particular features in experimentally-relevant noise models. One notable example of this is provided by Ref. [47], which constructs a very simple local Lindbladian decoder that corrects heralded erasure errors. While these will never be the sole source of errors, there are certainly scenarios in which they dominate, and it is natural to wonder to what extent our decoder can be optimized for such cases.
@@
This question will be addressed in upcoming work [27]. We will rigorously show that the offline decoders presented in this work, as well as the field-based decoders proposed in Ref. [14], do not have thresholds for real-time decoding. Fortunately, we will also demonstrate how to modify the message-passing architecture in a way which does guarantee the existence of a threshold, given the assumption that the noise acts only on the quantum system, and not on the data manipulated by the classical control infrastructure.
@@ 2 Preliminaries
In this preparatory section, we fix notation, introduce the general framework within which our decoders are defined, and discusses the types of noise models considered in this work. The discussion in this section will focus exclusively on synchronous decoders; the extension to asynchronous decoders which operate in continuous time case will be given later in Sec. 6.
@@ 2.1 Automaton decoders
We begin by fixing notation. Our decoders are defined for topological stabilizer codes $\mathcal{C}$ built from qubits living on the links of a lattice $\Lambda$, which we will usually take to be $\mathbb{Z}^{D}_{L}$ with periodic boundary conditions (all of our results also hold for open boundary conditions with minor modifications; the details are provided in App. B). $\Lambda_{l}$ will denote the set of links in $\Lambda$, and $B_{\mathbf{r}}^{(\mathsf{p})}(\Delta)$ will denote the ball of radius $\Delta$ in $\Lambda$ centered at a lattice point $\mathbf{r}$ defined with respect to the $\mathsf{p}$ norm (when the choice of $\mathsf{p}$ is not important, the superscript will be omitted). The stabilizers in $\mathcal{C}$ will be denoted by $g_{\mathbf{r}}$ (with $\Lambda$ defined such that each lattice point hosts a single stabilizer). A logical state will refer to a density matrix $\rho_{\sf log}$ such that $\rho_{\sf log}g_{\mathbf{r}}=g_{\mathbf{r}}\rho_{\sf log}=\rho_{\sf log}$ for all $\mathbf{r}$. While we will sometimes use more general notation, for simplicity we will ultimately always specify to the case of $\mathbb{Z}_{2}$ topological codes.
Our decoders will only focus on correcting logical qubits whose logical operators are string-like, since it is errors associated with these logical sectors whose correction is hardest to accomplish in a local / parallelized way. If $D$ is large enough so that some of the logical operators are supported on submanifolds of dimension $d>1$, these logical sectors can easily be corrected locally using a variant of Toom’s rule [16], and we will as such ignore such logical sectors in what follows.
@@
with $m_{\mathbf{r}}\in\{\pm 1\}$ a variable which will store the eigenvalue of $g_{\mathbf{r}}$. We will use boldface font to denote the collection of classical variables at all sites, so that $\mathbf{m}=(\sigma_{\mathbf{r}_{1}},\sigma_{\mathbf{r}_{2}},\dots)$ and likewise for ${\bm{\sigma}}$. If an operator $\mathcal{O}$ either commutes or anti-commutes with every $g_{\mathbf{r}}$, so that $g_{\mathbf{r}}\mathcal{O}=(-1)^{\sigma_{\mathcal{O},\mathbf{r}}}\mathcal{O}g_{\mathbf{r}}$, its syndrome ${\bm{\sigma}}_{\mathcal{O}}$ will be defined as the string ${\bm{\sigma}}_{\mathcal{O}}=(\sigma_{\mathcal{O},\mathbf{r}_{1}},\sigma_{\mathcal{O},\mathbf{r}_{2}},\dots)$. Finally, we will use the notation $P_{\mathbf{s}}=\bigotimes_{{\mathbf{l}}\in\Lambda_{l}}X^{s_{1,{\mathbf{l}}}}Z^{s_{2,{\mathbf{l}}}}$ to denote a Pauli operator labeled by the string $\mathbf{s}\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{|\Lambda_{l}|}$.
@@
A synchronous automaton decoder $\mathcal{D}=(\mathcal{A},\mathcal{F},\mathcal{C})$ is a discrete-time dynamics determined by an automaton rule $\mathcal{A}$, a set of feedback operators $\mathcal{F}$, and a topological stabilizer code $\mathcal{C}$. For an input state $\rho$ on $\mathcal{H}_{\sf qu}$, the decoder first measures all of the stabilizers of $\mathcal{C}$, producing the initial state $\rho_{0}=\sum_{\bm{\sigma}}\Pi_{\bm{\sigma}}\rho\Pi_{\bm{\sigma}}\otimes|\mathbf{m}_{0},{\bm{\sigma}}\rangle\langle\mathbf{m}_{0},{\bm{\sigma}}|_{\sf cl}$, where $\mathbf{m}_{0}$ is a particular initialization of the control variables and $\Pi_{\bm{\sigma}}=\bigotimes_{\mathbf{r}}(1+\sigma_{\mathbf{r}}g_{\mathbf{r}})/2$. At each time step $t$ after this initialization, the following operations in sequence:
@@
    The $\mathbf{m}$ variables in $\mathcal{H}_{\sf cl}$ are updated according to the automaton rule $\mathcal{A}$, which uses the values of $\mathbf{m},{\bm{\sigma}}$ to update the values of $\mathbf{m}$; we will use notation where $\mathcal{A}(\mathbf{m},{\bm{\sigma}})[m_{\mathbf{r}}]$ denotes the value that $m_{\mathbf{r}}$ takes upon applying $\mathcal{A}$ [^10] We can in general allow $\mathcal{A}$ to involve randomness, but for simplicity will stick with notation where $\mathcal{A}$ is deterministic. Formally, this is done using a channel $\Phi_{\mathcal{A}}$ with Kraus operators
@@
    $\mathcal{K}_{\mathbf{m},{\bm{\sigma}}}=\mathcal{F}(\mathbf{m},{\bm{\sigma}})\otimes|\mathbf{m},{\bm{\sigma}}{\bm{\sigma}}_{\mathcal{F}}\rangle\langle\mathbf{m},{\bm{\sigma}}|.$ (7)
@@ 2.2 Noise models
The input states our decoders will be tasked with decoding are obtained by subjecting logical states to certain types of noise channels. As mentioned in the introduction, we will restrict our attention to $p$-bounded stochastic noise channels, viz. channels which apply an incoherent mixture of Pauli operators to an input state, with the probability to apply any particular Pauli operator $P$ being exponentially suppressed in the size of $P$’s support. This is essentially the most general incoherent noise model one could consider, as no assumption on the spatial correlations between noise events are made. Importantly, this definition does not allow for any coherence between errors (the extension of the sparsity results in Sec. 4.1 to this case is somewhat involved, and will be deferred to future work [27]).
@@
We will only be interested in CSS codes in this work, and will only consider decoders $\mathcal{D}$ where the classical variables and syndromes can be separated as $\mathbf{m}=(\mathbf{m}^{X},\mathbf{m}^{Z})$, ${\bm{\sigma}}=({\bm{\sigma}}^{X},{\bm{\sigma}}^{Z})$, the automaton rule factored as $\mathcal{A}=\mathcal{A}^{X}\circ\mathcal{A}^{Z}$ with $\mathcal{A}^{X/Z}$ only depending on $\mathbf{m}^{X/Z},{\bm{\sigma}}^{Z/X}$, and the feedback operators split as $\mathcal{F}(\mathbf{m},{\bm{\sigma}})=\mathcal{F}^{X}(\mathbf{m}^{X},{\bm{\sigma}}^{Z})\mathcal{F}^{Z}(\mathbf{m}^{Z},{\bm{\sigma}}^{X})$, where $\mathcal{F}^{X/Z}$ are built entirely as tensor products of $X$ and $Z$ operators, respectively. We can accordingly split the decoder into $X$ and $Z$ parts as $\mathcal{D}=\mathcal{D}^{X}\circ\mathcal{D}^{Z}$, and as discussed in detail in Ref. [15], it suffices to separately consider the performance of $\mathcal{D}^{X/Z}$. While the $X$ and $Z$ errors will in general be correlated with one another, the generality of our $p$-bounded definition means that if we marginalize over e.g. the $Z$ errors, the resulting pattern of $X$ errors remains $p$-bounded. This means that if we construct $\mathcal{D}^{X/Z}$ to have thresholds under all $p$-bounded noise channels that apply only $X/Z$-type noise, $\mathcal{D}$ will have a threshold under general $p$-bounded noise channels. Using this fact, in the rest of this paper we will focus only on noise channels $\mathcal{E}$ which apply $p$-bounded $X$ noise, taking
@@
As a final comment, we note that while the $p$-bounded property is often referred to as implying that the noise has “decaying spatial correlations”, this is not true in any geometric sense: $\mathsf{P}({\mathbf{l}},{\mathbf{l}}^{\prime}\in\mathsf{N})$ needn’t factor as $\mathsf{P}({\mathbf{l}}\in\mathsf{N})\mathsf{P}({\mathbf{l}}^{\prime}\in\mathsf{N})$, even for arbitrarily large $||{\mathbf{l}}-{\mathbf{l}}^{\prime}||$. Rather, $p$-boundedness simply implies that the marginal probability of observing a particular pattern $A$ in the noise $\mathsf{N}$ is exponentially suppressed in the size of $A$. In the case where connected correlations between points in $\mathsf{N}$ decay exponentially in distance, we will say $\mathcal{E}$ has local correlations.
Figure 3: A schematic of the message passing architecture. a) The message-passing system in 1D. Anyons (green circles) live on a 1D chain, each site of which hosts two message fields $m^{\pm}$. Once the messages have equilibrated, the value of $m^{+}_{x}$ at site $x$ is equal to the distance between site $x$ and the nearest anyon to the left of site $x$; in the figure, a red number $n\rightarrow$ indicates that $m^{+}=n$ at this site. Similarly, the value of $m^{-}_{x}$ is determined by the distance between $x$ and the nearest anyon to the right of $x$; in the figure these are indicated by the blue numbers with $\leftarrow$ arrows. With the message fields shown in the figure, the left anyon will move right, and the right anyon will move left. b) The message-passing system in 2D. Each site now hosts four message fields $m^{\pm a}$, $a=x,y$, with the figure indicating the fields created by an anyon in the indicated location. $m^{\pm a}$ fields indicate the distance (in the $\infty$ norm) from an anyon in the $\mp{\mathbf{\hat{a}}}$ direction, and propagate along the ±$\pm a$ section of the anyon’s lightfront; in the figure, these lightfront segments are indicated by the shaded regions. The $m^{\pm a}$ field at a site $\mathbf{r}$ is updated by consulting the three sites at a infinity-norm distance of 1 in the $\mp{\mathbf{\hat{a}}}$ direction from $\mathbf{r}$. As an example, the site outlined in purple is updated according to the three sites indicated by the arrows. c) Movement of anyons in 2D. In the limit where messages are transmitted instantly, each anyon moves in the direction of its nearest neighbor (with distance calculated using the $\infty$ norm). If an anyon’s nearest neighbor is located in the ±$\pm a$ section of that anyon’s instantaneous lightfront, the anyon moves along ±$\pm{\mathbf{\hat{a}}}$. Thus the anyon $a$ in the figure moves along $+{\mathbf{\hat{x}}}$ towards anyon $b$, anyon $b$ moves along $+{\mathbf{\hat{y}}}$ towards anyon $c$, and so on. When the message transmission speed is finite, each anyon moves instead towards the nearest anyon in its past lightfront.

### v4

@@ 1.2 Comparison with prior work
The decoders most similar to ours—which served as direct inspiration for the present work—are the “field-based” CA surface code decoders of Refs. [13, 14]. These decoders transmit information about syndrome locations in a physically natural fashion, using a scalar field $\phi$ updated according to a lattice discretization of the Poisson equation
@@
For the purposes of this work, the most significant drawback of these decoders is that they do not meet our definition of locality, with the $D_{\phi}=3$ model requiring a $(\log L)^{2}$-divergent classical communication speed (in app. D, we argue that all PDE-based models suffer from this problem). Our decoder can thus be viewed as an improvement of this construction which brings down the communication speed to an $O(1)$ constant and reduces the number of control bits used per site from $\omega((\log L)^{3})$ [13] to $O(\log\log L)$.
@@
Aside from field-based decoders, the two existing types of cellular automaton decoders for the 2D toric code are Harrington’s decoder [12] and its extensions [40, 46] (see also [17, 20]), and the decoder of Ref. [15], which can be regarded as a quantum extension of Tsirelson’s automaton [47] (these decoders are designed for online decoding, but they can of course be used for offline decoding as well). Both of these decoders can loosely be thought of as eliminating anyons according to a locally-implemented RG scheme, with the different length scales involved in the RG process hard-coded into the structure of the dynamics. For the purposes of this work, drawbacks of these approaches are that they are only naturally defined on systems of size $n^{l}$ for integers $n,l$, and are numerically observed to have thresholds at rather low error rates, with the simplest formulation of Ref. [15] achieving $p_{c}\lesssim 2\%$ for offline decoding. These features make them—at least in their current form—less practical for experimental implementation and numerical study. [^8] In 2D, these decoders are also not naturally translation invariant in space or time, although this problem can be fixed formally by adding additional clock registers at each site and initializing the dynamics with an appropriate (fine-tuned) choice of initial conditions for the classical variables.
@@ 1.3 Outlook
While our explicit construction is specialized to the setting of topological codes defined on $\mathbb{Z}^{D}$, it should be straightforward to generalize the message passing architecture to codes defined on more general types of graphs. From an implementation perspective, it would also be valuable to quantitatively investigate different types of message passing schemes, and to systematically study how the numerical value of the threshold depends on the choice of architecture and parameters involved (such as the message transmission speed). Relatedly, it could be fruitful to study how decoding can be optimized to take advantage of particular features in experimentally-relevant noise models. One notable example of this is provided by Ref. [48], which constructs a very simple local Lindbladian decoder that corrects heralded erasure errors. While these will never be the sole source of errors, there are certainly scenarios in which they dominate, and it is natural to wonder to what extent our decoder can be optimized for such cases.
@@
This question is addressed in Ref. [27], where we rigorously show that the offline decoders presented in this work, as well as the field-based decoders proposed in Ref. [14], do not have thresholds for real-time decoding. Fortunately, that work also demonstrates how to modify the message-passing architecture in a way which does guarantee the existence of a threshold, at least given the assumption that the noise acts only on the quantum system, and not on the data manipulated by the classical control infrastructure.
@@ 2 Preliminaries
In this preparatory section, we fix notation, introduce the general framework within which our decoders are defined, and discusses the types of noise models considered in this work. To streamline the presentation, the discussion in this section will make two simplifications. First, we will focus exclusively on synchronous decoders that operate in discrete time. The extension to asynchronous Lindbladian decoders which operate in continuous time case will be given later in sec. 6. Second, we will explicitly adopt notation appropriate for the $D$-dimensional surface code. The generalization to arbitrary (potentially non-Abelian) topological codes is immediate, and will be deffered to sec. 5.4.
As a final preliminary comment, we note that we will focus solely on correcting logical qubits whose logical operators are string-like, since it is errors associated with these logical sectors whose correction is hardest to accomplish in a local / parallelized way. If $D$ is large enough so that some of the logical operators are supported on submanifolds of dimension $d>1$, these logical sectors can easily be corrected locally using a variant of Toom’s rule [16], and we will consequently ignore such logical sectors in what follows.
@@ 2.1 Automaton decoders
We begin by fixing notation. Our decoders are defined for codes built from qubits living on the links of a square lattice $\Lambda$, which we will usually take to be $\mathbb{Z}^{D}_{L}$ with periodic boundary conditions (all of our results also hold for open boundary conditions with minor modifications; the details are provided in app. B). $\Lambda_{l}$ will denote the set of links in $\Lambda$, and $B_{\mathbf{r}}^{(\mathsf{p})}(\Delta)$ will denote the ball of radius $\Delta$ in $\Lambda$ centered at a lattice point $\mathbf{r}$ defined with respect to the $\mathsf{p}$ norm (when the choice of $\mathsf{p}$ is not important, the superscript will be omitted). The stabilizers which define the code space will be denoted by $g_{\mathbf{r}}$, with $\Lambda$ defined such that each lattice point hosts a single stabilizer. A logical state will refer to a density matrix $\rho_{\sf log}$ such that $\rho_{\sf log}g_{\mathbf{r}}=g_{\mathbf{r}}\rho_{\sf log}=\rho_{\sf log}$ for all $\mathbf{r}$.
@@
with $\sigma_{\mathbf{r}}\in\{\pm 1\}$ a variable which will store the eigenvalue of $g_{\mathbf{r}}$. We will use boldface font to denote the collection of classical variables at all sites, so that $\mathbf{m}=(m_{\mathbf{r}_{1}},m_{\mathbf{r}_{2}},\dots)$ and likewise for $\bm{\sigma}$}. If an operator $\mathcal{O}$ either commutes or anti-commutes with every $g_{\mathbf{r}}$, so that $g_{\mathbf{r}}\mathcal{O}=(-1)^{\sigma_{\mathcal{O},\mathbf{r}}}\mathcal{O}g_{\mathbf{r}}$, its syndrome ${\bm{\sigma}}_{\mathcal{O}}$ will be defined as the string ${\bm{\sigma}}_{\mathcal{O}}=(\sigma_{\mathcal{O},\mathbf{r}_{1}},\sigma_{\mathcal{O},\mathbf{r}_{2}},\dots)$. Finally, we will use the notation $P_{\mathbf{s}}=\bigotimes_{{\mathbf{l}}\in\Lambda_{l}}X^{s_{1,{\mathbf{l}}}}Z^{s_{2,{\mathbf{l}}}}$ to denote a Pauli operator labeled by the string $\mathbf{s}\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{|\Lambda_{l}|}$.
@@
A synchronous automaton decoder $\mathcal{D}=(\mathcal{A},\mathcal{F})$ is a discrete-time dynamics determined by an automaton rule $\mathcal{A}$ and a set of feedback operators $\mathcal{F}$. For an input state $\rho$ on $\mathcal{H}_{\sf qu}$, the decoder first measures the topological charges at each lattice site, producing the initial state $\rho_{0}=\sum_{\bm{\sigma}}\Pi_{\bm{\sigma}}\rho\Pi_{\bm{\sigma}}\otimes|\mathbf{m}_{0},{\bm{\sigma}}\rangle\langle\mathbf{m}_{0},{\bm{\sigma}}|_{\sf cl}$, where $\mathbf{m}_{0}$ is a particular initialization of the control variables and $\Pi_{\bm{\sigma}}=\bigotimes_{\mathbf{r}}(1+\sigma_{\mathbf{r}}g_{\mathbf{r}})/2$. At each time step $t$ after this initialization, the following operations in sequence:
@@
    The $\mathbf{m}$ variables in $\mathcal{H}_{\sf cl}$ are simultaneously updated according to the automaton rule $\mathcal{A}$, which uses the values of $\mathbf{m},{\bm{\sigma}}$ to update the values of $\mathbf{m}$; we will use notation where $\mathcal{A}(\mathbf{m},{\bm{\sigma}})[m_{\mathbf{r}}]$ denotes the value that $m_{\mathbf{r}}$ takes upon applying $\mathcal{A}$. [^9] We can in general allow $\mathcal{A}$ to involve randomness, but for simplicity will stick with notation where $\mathcal{A}$ is deterministic. Formally, this is done using a channel $\Phi_{\mathcal{A}}$ with Kraus operators
@@
    $\mathcal{K}_{\mathbf{m},{\bm{\sigma}}}=\mathcal{F}(\mathbf{m},{\bm{\sigma}})\otimes|\mathbf{m},{\bm{\sigma}}{\bm{\sigma}}_{\mathcal{F}}\rangle\langle\mathbf{m},{\bm{\sigma}}|,$ (7)
    where ${\bm{\sigma}}{\bm{\sigma}}_{\mathcal{F}}$ denotes element-wise multiplication.
@@
The word “synchronous” is used here to refer to the fact that the automaton updates and feedback are applied synchronously, with all sites in the system updating in union at each time step.
@@ 2.2 Noise models
The input states our decoders will be tasked with decoding are obtained by subjecting logical states to certain types of noise channels. As mentioned in the introduction, we will restrict our attention to $p$-bounded stochastic noise channels, viz. channels which apply an incoherent mixture of Pauli operators to an input state, with the probability to apply any particular Pauli operator $P$ being exponentially suppressed in the size of $P$’s support. This is essentially the most general incoherent noise model one could consider, as no assumption on the spatial correlations between noise events are made. Importantly, this definition does not allow for any coherence between errors, although this assumption can be removed following the techniques of [34, 31].
@@
For CSS codes like the surface code—which is the main example studied in this work—we will always consider decoders $\mathcal{D}$ where the classical variables and syndromes can be separated as $\mathbf{m}=(\mathbf{m}^{X},\mathbf{m}^{Z})$, ${\bm{\sigma}}=({\bm{\sigma}}^{X},{\bm{\sigma}}^{Z})$, the automaton rule factored as $\mathcal{A}=\mathcal{A}^{X}\circ\mathcal{A}^{Z}$ with $\mathcal{A}^{X/Z}$ only depending on $\mathbf{m}^{X/Z},{\bm{\sigma}}^{Z/X}$, and the feedback operators split as $\mathcal{F}(\mathbf{m},{\bm{\sigma}})=\mathcal{F}^{X}(\mathbf{m}^{X},{\bm{\sigma}}^{Z})\mathcal{F}^{Z}(\mathbf{m}^{Z},{\bm{\sigma}}^{X})$, where $\mathcal{F}^{X/Z}$ are built entirely as tensor products of $X$ and $Z$ operators, respectively. We can accordingly split the decoder into $X$ and $Z$ parts as $\mathcal{D}=\mathcal{D}^{X}\circ\mathcal{D}^{Z}$, and as discussed in detail in Ref. [15], it suffices to separately consider the performance of $\mathcal{D}^{X/Z}$. While the $X$ and $Z$ errors will in general be correlated with one another, the generality of our $p$-bounded definition means that if we marginalize over e.g. the $Z$ errors, the resulting pattern of $X$ errors remains $p$-bounded. This means that if we construct $\mathcal{D}^{X/Z}$ to have thresholds under all $p$-bounded noise channels that apply only $X/Z$-type noise, $\mathcal{D}$ will have a threshold under general $p$-bounded noise channels. Using this fact, until we discuss general topological codes in sec. 5.4 we will focus only on noise channels $\mathcal{E}$ which apply $p$-bounded $X$ noise, taking
@@
As a final comment, we note that while the $p$-bounded property is often referred to as implying that the noise is “local” and has “decaying spatial correlations”, this is not true in any geometric sense: $\mathsf{P}({\mathbf{l}},{\mathbf{l}}^{\prime}\in\mathsf{N})$ needn’t factor as $\mathsf{P}({\mathbf{l}}\in\mathsf{N})\mathsf{P}({\mathbf{l}}^{\prime}\in\mathsf{N})$, even for arbitrarily large ${\mathbf{l}}-{\mathbf{l}}^{\prime}||$. Rather, $p$-boundedness simply implies that the marginal probability of observing a particular pattern $A$ in the noise $\mathsf{N}$ is exponentially suppressed in the size of $A$. In the case where connected correlations between points in $\mathsf{N}$ decay exponentially in distance, we will say $\mathcal{E}$ has (spatially) local correlations.
Figure 3: A schematic of the message passing architecture. a) The message-passing system in 1D. Anyons (green circles) live on a 1D chain, each site of which hosts two message fields $m^{\pm}$. Once the messages have equilibrated, the value of $m^{+}_{x}$ at site $x$ is equal to the distance between site $x$ and the nearest anyon to the left of site $x$; in the figure, a red number $n\rightarrow$ indicates that $m^{+}=n$ at this site. Similarly, the value of $m^{-}_{x}$ is determined by the distance between $x$ and the nearest anyon to the right of $x$; in the figure these are indicated by the blue numbers with $\leftarrow$ arrows. With the message fields shown in the figure, the left anyon will move right, and the right anyon will move left. b) The message-passing system in 2D. Each site now hosts four message fields $m^{\pm a}$, $a=x,y$, with the figure indicating the fields created by an anyon in the indicated location. $m^{\pm a}$ fields indicate the distance (in the $\infty$ norm) from an anyon in the $\mp{\mathbf{\hat{a}}}$ direction, and propagate along the $\pm a$ section of the anyon’s lightfront; in the figure, these lightfront segments are indicated by the shaded regions. The $m^{\pm a}$ field at a site $\mathbf{r}$ is updated by consulting the three sites at a infinity-norm distance of 1 in the $\mp{\mathbf{\hat{a}}}$ direction from $\mathbf{r}$. As an example, the site outlined in purple is updated according to the three sites indicated by the arrows. c) Movement of anyons in 2D. In the limit where messages are transmitted instantly, each anyon moves in the direction of its nearest neighbor (with distance calculated using the $\infty$ norm). If an anyon’s nearest neighbor is located in the $\pm a$ section of that anyon’s instantaneous lightfront, the anyon moves along $\pm{\mathbf{\hat{a}}}$. Thus the anyon $a$ in the figure moves along $\mathbf{\hat{x}}$} towards anyon $b$, anyon $b$ moves along $\mathbf{\hat{y}}$} towards anyon $c$, and so on. When the message transmission speed is finite, each anyon moves instead towards the nearest anyon in its past lightfront.